OpenStudy (anonymous):

Assume that a population of ticks is growing exponentially. Find the number of ticks after 25 days if an initial population of 100 ticks triples in 10 days.

OpenStudy (australopithecus):

A=Pe^(rt)

OpenStudy (australopithecus):

im pretty sure is the formula

OpenStudy (anonymous):

100000

OpenStudy (australopithecus):

First you need to figure out the rate we know that A = 300 P = 100 t = 10 r = ? Once we figure out the rate we can figure out how many ticks we have in 25 days

OpenStudy (anonymous):

\[100\times 3^{\frac{25}{10}}\] will work

OpenStudy (australopithecus):

Although satellites method may be easier

OpenStudy (anonymous):

no need to find the rate, you have it. it triples every 10 days

OpenStudy (australopithecus):

My method is longer but it still works satellite :)

OpenStudy (australopithecus):

although your method is smarter

OpenStudy (anonymous):

otherwise you have to solve an exponential equation, then go back and replace t by 25. too much work for me, and also less accurate since you will probably round the rate

OpenStudy (anonymous):

where is 300 from on a?

OpenStudy (anonymous):

yes of course it works.

OpenStudy (australopithecus):

depends how you perform the calculaton satellite :\ so the accuracy is not a problem

OpenStudy (australopithecus):

but yes your method is better no need to rub it in :)

OpenStudy (ash2326):

Population (P) grows exponentially, given that it triples in 10 days, We can write as \[P=p \times 3^{t/10}\] p = initial population p=100 \[P=100\times 3^{t/10}\] After 25 days P will be \[P=100\times 3^{25/10}\]

OpenStudy (anonymous):

not better, just easier

OpenStudy (anonymous):

so the answer i got is 1558.84 is that correct?

OpenStudy (anonymous):

?